Pixel:

Two cones are similar. The volume of the larger cone is 27 cm³ and the volume of the smaller cone is 8 cm³. The height of the smaller cone is 2 cm. What is the height of the larger cone?

Hanna:

Volume of cone, V = (1/3)πr^2h let H, h and R and r are height and radius of larger and smaller cones respectively (1/3)πR^2 H /(1/3)πr^2 h = (27/8) => R^2 H / r^2 h = (27/8)


eqn(1) since cones are similar R/r = H/h ==> R^2 /r^2 = H^2 /h^2 substitute in eqn (1) H^3 /h^3 = (27/8) H/h = 3/2 since h = 2 cm H/2 = 3/2 => H = 3 cm Hope that helps. :)